AirWizz Products, a manufacturer of aircraft landing gear, makes 1,300 units each year of a special valve used in assembling one of its products. The unit cost of producing this valve includes variable costs of $73 and fixed costs of $60. The valves could be purchased from an outside supplier at $80 each. If the valve were purchased from the outside supplier, 40% of the total fixed costs incurred in producing this valve could be eliminated. Buying the valves from the outside supplier instead of making them would cause the company's operating income to:
Answer #1

Differential analysis

Make Buy
Variable cost 1300*73 = 94900
Fixed cost 1300*60*40% = 31200
Purchase cost 1300*80 = 104000
Total relevant cost 126100 104000

Buying the valves from the outside supplier instead of making them would cause the company's operating income to:

Increase by (126100-104000) = 22100